OpenStudy (anonymous):

The length of a rectangle yard is 30 meters the perimeter Is at 90 meters

OpenStudy (anonymous):

perimeter = 2 * length + 2 * width at most 90 m = 2*30 m + 2 * width at most 90 m = 60 m + 2 * width at most 30 m = 2 * width at most 15 m = width

OpenStudy (anonymous):

Btw, I love fans & medals. :)

OpenStudy (anonymous):

I can tell ^^ It is better to show work for the user instead of just giving an answer. @connexuz.com I keep running into you

OpenStudy (anonymous):

I did show my work. >_>